40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 15650     Entry time: Thu Oct 29 09:50:12 2020
Author: anchal 
Type: Summary 
Category: Calibration 
Subject: Preliminary calibration measurement taken 

I went to 40m yesterday at around 2:30 pm and Koji showed me how to acquire lock in different arms and for different lasers. Finally, we took a preliminary measurement of shaking the ETMX at some discrete frequencies and looking at the beatnote frequency spectrum of X-end laser's fiber-coupled IR and Main laser's IR pick-off.


Basic controls and measurement 101 at 40m

  • I learned a few things from Koji about how to align the cavity mirrors for green laser or IR laser.
  • I learned how to use ASS and how to align the green end laser to the cavity. I also found out about the window at ETMX chamber where we can directly see the cavity mode, cool stuff.
  • Koji also showed me around on how to use diaggui and awggui for taking measurements with any of the channels.

Preliminary measurement for calibration scheme

We verified that we can send discrete frequency excitation signals to ETMX actuators directly and see a corresponding peak in the spectrum of beatnote frequency between fiber-coupled X-end IR laser and main laser IR pickoff.

  • I sent excitation signal at 200 Hz, 250 Hz and 270 Hz at C1:SUS-ETMX_LSC_EXC channel using awggui with an amplitude of 100 cts and gain of 2.
  • I measured corresponding peaks in the beatnote spectrum using diaggui.
  • Page 1 shows the ASD data for the 4 measurements taken with Hanning window and averaging of 10.
  • Page 2 shows close up Spectrum data for the 4 measurements taken with flattop window and averaging of 10.
  • I converted this frequency signal into displacement by using conversion factor \nu_{FSR}/\frac{\lambda}{2} or \frac{L \lambda}{c}.

If full interferometer had been locked, we could have used the DARM error signal output to calibrate it against this measurement.

Data

Attachment 1: PreliminaryCalibrationData.pdf  178 kB  Uploaded Thu Oct 29 11:25:27 2020  | Hide | Hide all
PreliminaryCalibrationData.pdf PreliminaryCalibrationData.pdf
ELOG V3.1.3-